OpenStudy (anonymous):

[7.06] Choose the correct simplification of 8x2(5x + 2x2 − 3). 16x4 − 40x3 + 24x2 16x4 + 40x3 − 24x2 40x4 + 16x3 − 5x2 40x4 − 10x3 + 5x2

OpenStudy (anonymous):

\[8x^2(5x + 2x^2 − 3)\] ??

OpenStudy (anonymous):

Distribute the \(8x^2 \) to all the terms inside the \(( ) \)

OpenStudy (anonymous):

will it be this 16x4 − 40x3 + 24x2

OpenStudy (anonymous):

Fellica watch the minus sign.

OpenStudy (anonymous):

so this one 16x4 + 40x3 − 24x2

OpenStudy (anonymous):

So: 8x^2 * 5x = 40x^3 8x^2 * 2x^2 = 16x^4 8x^2 * -3 = -24x^2 then: 16x^4 + 40x^3 - 24x^2

OpenStudy (anonymous):

yup good :) @Felicia123

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

[7.07] Which expression is equivalent to the area of metal sheet required to make this square shaped traffic sign? A square shaped traffic sign is shown with length of one side labeled as 2x − 1. 4x2 - 1 4x2 + 1 4x2 + 4x - 1 4x2 - 4x + 1

OpenStudy (anonymous):

link broken

OpenStudy (anonymous):

Cookie Error: The system is unable to access the cookie that was assigned to your computer upon login. The most likely cause of this problem is an improper access to the course via a bookmark, browser window that was left open, or an improper exit from the system. To resolve the problem you will need to: Log into your course Delete your browser cookies Click the "Logoff" button Reboot your computer Sign back into your account Make sure you properly logoff and close your browser window each time you exit the system If you need assistance resolving this problem please follow the steps below: Click on the "Technical Support" link within your course. Click the "Educator Help Desk" sublink, located just below the "Technical Support" button you just clicked. Select that the "Nature of your problem" is with the "Educator System". This will route the request to our team. In the problem description field, provide a brief description of the problem you are experiencing. In this case, something along the lines of "When I click on Lessons I get a cookie error" will be more than sufficient. When you click the "Submit" button this will send your problem description to our team along with your account information and some information about the browser you are using. We will then reset your account as quickly as possible and provide you with the appropriate steps to reset your browser on your computer.

OpenStudy (anonymous):

tat is what it says in the link ^

OpenStudy (anonymous):

ok hold on lol

OpenStudy (anonymous):

Are you looking for the area?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

help me please i don't know what to do and i don't know what i am looking for help

OpenStudy (anonymous):

Well It looks like a square so for area do \[(2x-1)^{2} = (2x-1) \times (2x-1)\] If its perimeter do. \[P = 4(2x -1)\]

OpenStudy (anonymous):

be this one 4x2 + 4x - 1 mebs

OpenStudy (anonymous):

or this 4x2 + 1

OpenStudy (anonymous):

No it would bee. 4x2 -4x +1

OpenStudy (anonymous):

got to be this one 4x2 + 1

OpenStudy (anonymous):

why?